# Non-Markovian Reduced Models to Unravel Transitions in Non-equilibrium Systems

### v1

@@ V.2 Non-Markovian optimal reduced models
Recall that our general goal is the derivation of reduced models able to capture the emergence of noise-driven spatiotemporal patterns and predict their transitions triggered by the subtle coupling between noise and the nonlinear dynamics.
@@
In other words, the reduced state space is spanned by low modes that are not forced stochastically; i.e. $q=m_{c}$ in (77). The subspace $H_{\mathfrak{s}}$ is taken to be the orthogonal complement of $H_{\mathfrak{c}}$ in $L^{2}(0,L)$ and contains the stochastically forced modes; see (77).
@@
$\Phi_{n}$ $(\tau_{n},\bm{\lambda},X,t;\omega)$ (80)
$=\sigma_{n}\big{(}W^{n}_{t}(\omega)-e^{\tau\lambda_{n}}W^{n}_{t-\tau}(\omega)\big{)}+Z^{n}_{\tau}({\bm{\lambda}},t;\omega)$
$\qquad+\sum_{i,j,k=1}^{q}\Big{(}E_{ijk}^{n}(\tau,\bm{\lambda})C_{ijk}^{n}X_{i}X_{j}X_{k}\Big{)},$
for each $n$ in $(q+1,N)$. Here $E_{ijk}^{n}$ is given in (44), and $Z^{n}_{\tau}({\bm{\lambda}},t;\omega)$ is given by (41).
@@
The optimal reduced system (82) can be further expanded as
@@
$+\sum_{n,n^{\prime},n^{\prime\prime}=q+1}^{N}C_{nn^{\prime}n^{\prime\prime}}^{i}\Phi_{n}(\tau^{*}_{n},\bm{y},t;\omega)\Phi_{n^{\prime}}(\tau^{*}_{n^{\prime}},\bm{y},t;\omega)\Phi_{n^{\prime\prime}}(\tau^{*}_{n^{\prime\prime}},\bm{y},t;\omega),$
in which we have dropped the dependence on $\bm{\lambda}$ of the $\Phi_{n}$, to simplify the notations. Here also, the coefficients $C_{ijk}^{\ell}$ are those defined in (81) for $1\leq i,j,k,\ell\leq N$. Note that the required $Z^{n}$-terms to simulate Eq. (83) are advanced in time by solving the corresponding Eq. (47). A total of $N-q$ such equations are solved, each corresponding to a single random coefficient per mode to parameterize. Thus, a total of, $N=q+(N-q)$, ODEs with random coefficients are solved to obtain a reduced model of the amplitudes of the $q$ first modes. We call Eq. (83) together with its auxiliary equations (47), the optimal reduced model for the sACE.
@@
Noteworthy is that an approximation of the sACE solution paths can be obtained from the solutions $\bm{y}(t,\omega)$ to the optimal reduced model. Such an approximation consists of simply lifting the surrogate low-mode amplitude $\bm{y}(t,\omega)$ simulated from Eq. (84), to $H_{\mathfrak{s}}$ via the optimal parameterization, $\Phi_{{\bm{\tau}}^{\ast}}$, defined in Eq. (84). This approximation takes then the form
@@
Figure 4: Ginzburg-Landau free energy functional. Here, the Ginzburg-Landau free energy functional $\mathcal{E}(u)$ given by (75) is shown over the reduced state made of the two dominant modes’ amplitudes. A few solution paths navigating across this energy’s landscape (over $[0,T]$) are shown. The “shallower” local minima correspond to $\phi_{2}^{a/b}$ (sign-change profile), while the deeper ones correspond to ϕ1±superscriptsubscriptitalic-$\phi_{1}^{\pm}$ (constant-sign profile). The insets show examples of metastable states reached at $t=T$ by the sACE near these steady states. The rare paths are those trapped in the deeper wells due to the degenerate noise employed here; see Text.
Figure 5: Non-Markovian terms and rare path’s small-scale trigger. Panel A shows the small-scale features of the solution obtained from the full equation (sACE) responsible for triggering a rare transition path. The non-Markovian field, $\xi_{t}$ (defined by Eq. (LABEL:Non-Markov_field)), dominates the optimal parameterization, especially over the time interval $[0,10]$ (since $u_{\mathfrak{c}}\approx 0$ for this interval). Panel B shows that it provides an accurate parameterization of these small-scale features. Panels C to F show how the optimal reduced model effectively captures the dynamics arising from the interplay between the non-linear terms and noise in the system.
Figure 6: Non-Markovian terms and typical path’s small-scale trigger. Same as Fig. 5 but for a typical path.
@@ V.3 Why energy flows in a certain direction: The role of non-Markovian effects
Determining why energy primarily accumulates in the $\bm{e}_{2}$ direction is not straightforward. It relates to how noise initially affecting unresolved modes (modes $5$ to $8$) propagates through the system’s nonlinear terms over time, gradually influencing the larger-scale resolved modes. Since the initial state ($u0=0$) has no energy, this noise transmission to larger scales takes some time. Our experiments show that during the interval $(0,10)$, the solution is dominated by these small-scale features inherited from the noise forcing (see Table 2 and Fig. 3).
This finding has important implications for closure methods: to accurately capture either rare or typical system’s transition paths, the underlying parameterization must effectively represent the system’s small-scale response to the stochastic forcing near the beginning ($t=0$).
Figures 5 and 6 illustrate how well the optimal parameterization and reduced model capture the subtle differences in the small-scale triggers that ultimately lead to either rare or typical events in the full sACE system. The non-Markovian terms are central in the good capture of these small-scale triggers. Indeed, as the sACE low-mode’s amplitude is close to zero over $(0,10)$ (see Table 2) the optimal parameterization (Eq. (84)) is dominated over this interval by the (optimized) non-Markovian field
@@
$\hskip 142.26378pt+Z^{n}_{\tau_{n}^{\ast}}({\bm{\lambda}},t;\omega),$
before the nonlinear effects kick in.
@@
After, this transient dynamics, when energy gets pumped into the low modes, the role of the nonlinear terms in the stochastic parameterization (80) becomes key for capturing the sACE dynamics’ average behavior, as emphasized in Section V.5 below.
@@ V.4 Rare and typical events statistics from non-Markovian reduced models
Thus, the optimal reduced system accurately reproduces the probability distribution of both common (frequently occurring sign-change profiles) and rare (constant-sign profiles) final states at time $T$ (Fig. 7A). This is particularly impressive considering the model was trained on just a single path, yet generalizes well to predict the behavior across a large ensemble of out-of-sample paths.
Figure 8A shows the training stage over a typical transition path. The optimization of the free parameter $\tau$ in the stochastic paramterization $\Phi_{n}$ (see Eq. (80)) is executed over a typical transition path, by minimizing the normalized parameterization defects $Q_{n}(\tau,T)=\mathcal{Q}_{n}(\tau,T)/\overline{\big{|}u_{n}(t)\big{|}^{2}}$ ($\mathcal{Q}_{n}(\tau,T)$ defined in Eq. (51)) for all the relevant $n$ (here $n=5,\cdots,8$). We observe that $Q_{5}$ and $Q_{6}$ exhibit clear minima, whereas the other modes have their parameterization defect not substantially diminished at finite $\tau$-values compared to its asymptotic value (as $\tau\rightarrow\infty$). Same features are observed over a rare transition path; see Figure 8B.
Recall that our reduced system is an ODE system whose stochasticity comes exclusively from its random coefficients as the resolved modes in the full system are not directly forced by noise. It is this specificity that enables the reduced equations to effectively predict noise-induced transitions, including rare events. This remarkable capability relies on the optimized memory content embedded within the path-dependent, non-Markovian coefficients.
Thus, the optimized, non-Markovian and nonlinear reduced equation (83) is able to track efficiently how the noise, acting on the “unseen” part of the system, interacts with the resolved modes through these non-Markovian coefficients. While these coefficients effectively capture the fluctuations for a good reduction of the sACE system, they alone are insufficient to handle regimes with weak time-scale separation such as dealt with here. Understanding the average behavior of the resolved dynamics is equally important. The next section explores how our parameterization, particularly its nonlinear terms, play a crucial role in capturing this average behavior.
@@ V.5 Weak time-scale separation, nonlinear terms and conditional expectation
We assess the ability of our reduced system to capture this average dynamical behavior in the physical space. In that respect, figures 7B and 7D show the expected profiles of the sACE’s states obtained at $t=T$ (solid black curves) from the sACE (74), along with their standard deviation (grey shaded areas), for the ensemble used for Fig. 7A. Similarly, Figures 7C and 7E show the expected profiles (solid red curves) with their standard deviation as obtained at $t=T$ from $u^{\mathrm{opm}}(T,x,\cdot)$ defined in Eq. (85), once the low-mode amplitude is simulated by the optimal reduced model (84) and lifted to $H_{\mathfrak{s}}$ by using the stochastic parameterization $\Phi_{{\bm{\tau}}^{\ast}}$.
@@
Figure 9: Markovian optimal parameterization for the 5th and 6th modes’ amplitudes. Panel A: The black curve corresponds to the 6th mode’s amplitude $u_{6}(t)=\langle u(t),\bm{e}_{6}\rangle$ with $u$ solving Eq. (74), against its first and second modes’ amplitudes, $u_{1}(t)$ and $u_{2}(t)$, while the others in $H_{\mathfrak{c}}$, $u_{3}(t)$ and $u_{4}(t)$, are set to their most probable values $p_{3}$ and $p_{4}$, respectively. The manifold shown here is obtained as graph of the Markovian parameterization, $(X_{1},X_{2})\mapsto\mathbb{E}[\Phi_{6}(\tau_{6}^{\ast},\bm{X},t;\cdot)]$, for $\bm{X}=(X_{1},X_{2},X_{3},X_{4})$ with $X_{3}=p_{3}$ and $X_{4}=p_{4}$ in (80). Panel B: Same for 5th mode’s amplitude $u_{5}(t)=\langle u(t),\bm{e}_{5}\rangle$ and $\mathbb{E}[\Phi_{5}(\tau_{5}^{\ast},\bm{X},t;\cdot)]$.
@@
Thus, from large-ensemble (online) simulations of our reduced model (Eq. (84)), we are able to accurately capture the average motion and second-order statistics of the original sACE. This success hinges on our parameterization’s ability to capture the average behavior (for the probability measure $\mu$ defined in (56)) of the unresolved variables during the offline training stage. Figure 9 demonstrates this for the fifth and sixth modes’ amplitudes of the sACE system. It highlights another key advantage of our parameterization optimization: enabling accurate prediction of the average motion. Notably, the nonlinear shape of the Markovian optimal parameterization (obtained after taking the expectation, see Figure 9) reflects the importance of the nonlinear terms in our stochastic parameterization (Eq. (80)) for achieving this accurate capture.
@@
Without the optimization stage, our parameterizations can encounter problems due to small values that the spectral gaps $\delta_{ijk}^{n}$ can take in denominators. As a reminder, small spectral gaps are indeed a known limitation of traditional invariant manifold techniques, often leading to inaccurate results in this case. Optimizing the parameter $\tau$ helps address this issue. It introduces corrective terms like $(1-\exp(-\delta_{ijk}^{n})\tau)$ in the cubic coefficients ($E_{ijk}^{n}$, Eq. (44)) that effectively compensate for these small gaps. This is especially important for fifth and sixth (unresolved) modes whose parameterization defects show a clear minimum during optimization (Figure 8).
@@
Thus, the results in Figure 7 demonstrate our reduction approach’s effectiveness in handling weak time-scale separation scenarios, while also suggesting potential for further improvement (see Section VII). Our data-driven optimization of analytical parameterization formulas offers a compelling alternative to traditional multiscale methods. These methods, while providing rigorous error estimates, often require strong time-scale separation for the derivation of their amplitude equations. Additionally, they can achieve higher-order nonlinearities in their reduced equations too but through a different route: the Itô formula is applied after exploiting a slow timescale of the system (e.g., [49, Eq. (12)] and [56, Eq. (15)]). Integrating optimization concepts into this framework could be a promising avenue to explore how it performs in weak time-scale separation regimes.

### v2

@@ V.2 Non-Markovian optimal reduced models
Recall that our general goal is the derivation of reduced models able to reproduce the emergence of noise-driven spatiotemporal patterns and predict their transitions triggered by the subtle coupling between noise and the nonlinear dynamics.
@@
In other words, the reduced state space is spanned by low modes that are not forced stochastically; i.e. $q=m_{c}$ in (79). The subspace $H_{\mathfrak{f}}$ is taken to be the orthogonal complement of $H_{\mathfrak{c}}$ in $L^{2}(0,L)$ and contains the stochastically forced modes; see (79).
@@
$\Phi_{n}$ $(\tau,X,t;\omega)$ (82)
$=\sigma_{n}\big{(}W^{n}_{t}(\omega)-e^{\tau\lambda_{n}}W^{n}_{t-\tau}(\omega)\big{)}+Z^{n}_{\tau}(t;\omega)$
$\qquad+\sum_{i,j,k=1}^{q}\Big{(}E_{ijk}^{n}(\tau)C_{ijk}^{n}X_{i}X_{j}X_{k}\Big{)},$
for each $n$ in $(q+1,N)$. Here $E_{ijk}^{n}$ is given in (44), and $Z^{n}_{\tau}(t;\omega)$ is given by (41).
@@
The optimal reduced system (84), also referred to as the OPM closure hereafter, can be further expanded as
@@
$+\sum_{n,n^{\prime},n^{\prime\prime}=q+1}^{N}C_{nn^{\prime}n^{\prime\prime}}^{i}\Phi_{n}(\tau^{*}_{n},\bm{y},t;\omega)\Phi_{n^{\prime}}(\tau^{*}_{n^{\prime}},\bm{y},t;\omega)\Phi_{n^{\prime\prime}}(\tau^{*}_{n^{\prime\prime}},\bm{y},t;\omega).$
Here also, the coefficients $C_{ijk}^{\ell}$ are those defined in (83) for $1\leq i,j,k,\ell\leq N$. Note that the required $Z^{n}$-terms to simulate Eq. (85) are advanced in time by solving the corresponding Eq. (49). A total of $N-q$ such equations are solved, each corresponding to a single random coefficient per mode to parameterize. Thus, a total of, $N=q+(N-q)$, ODEs with random coefficients are solved to obtain a reduced model of the amplitudes of the $q$ first modes. We call Eq. (85) together with its auxiliary equations (49), the optimal reduced model for the sACE.
@@
Noteworthy is that an approximation of the sACE solution paths can be obtained from the solutions $\bm{y}(t,\omega)$ to the optimal reduced model. Such an approximation consists of simply lifting the surrogate low-mode amplitude $\bm{y}(t,\omega)$ simulated from Eq. (86), to $H_{\mathfrak{f}}$ via the optimal parameterization, $\Phi_{{\bm{\tau}}^{\ast}}$, defined in Eq. (86). This approximation takes then the form
@@
Figure 4: Ginzburg-Landau free energy functional. Here, the Ginzburg-Landau free energy functional $\mathcal{E}(u)$ given by (77) is shown over the reduced state made of the two dominant modes’ amplitudes. A few solution paths navigating across this energy’s landscape (over $[0,T]$) are shown. The “shallower” local minima correspond to $\phi_{2}^{a/b}$ (sign-change profile), while the deeper ones correspond to ϕ1±superscriptsubscriptitalic-$\phi_{1}^{\pm}$ (constant-sign profile). The insets show examples of metastable states, reached at $t=T$ by the sACE, located near these steady states. The rare paths are those trapped in the deeper wells due to the degenerate noise employed here; see Text.
Figure 5: Non-Markovian terms and rare path’s small-scale component. Panel A shows the small-scale component of the solution to the sACE, experiencing a rare transition path. Panel B shows the OPM parameterization of this small-scale component over the interval (0,10), prior the transition takes place. Panels C to F show the ability of the optimal reduced model, Eq. (85), to predict the large fluctuations experienced by the large-scale modes’ amplitudes, beyond this interval.
Figure 6: Non-Markovian terms and typical path’s small-scale component. Same as Fig. 5 but for a typical path.
@@ V.3 Why energy flows in a certain direction: The role of non-Markovian effects
Understanding why energy primarily accumulates in the $\bm{e}_{2}$ direction is complex. It stems from how noise initially affecting unresolved modes (modes $5$ to $8$) propagates through nonlinear interactions, gradually influencing larger-scale resolved modes. As the initial state ($u_{0}=0$) lacks energy, this noise-induced energy transfer to larger scales takes some time. Our analysis reveals that during the interval $I=(0,10)$, the solution is dominated by these small-scale features inherited from the noise forcing (Table 2 and Fig. 3).
This finding has significant implications for closure methods: to accurately reproduce both rare and typical transitions, the underlying parameterization must effectively represent the system’s small-scale response to stochastic forcing before the transition occurs.
Figures 5 and 6 illustrate how the optimal parameterization and reduced model (Eq. (86)) capture the subtle differences in the small-scale component that ultimately lead to rare or typical events in the full sACE system. Notably, for both rare (Figure 5) and typical paths (Figure 6), the low-mode amplitudes of the sACE solution are near zero during the interval $I=(0,10)$ (Panels C-F and Table 2). Therefore, it is the spatial patterns of the small-scale component, rather than its energy content, that control the triggering of rare or typical events (Figure 3A and 3D). The successful parameterization of these small-scale features (Figures 5B and 6B) is crucial for the optimal reduced model’s ability to reproduce the key behaviors of the sACE’s large-scale dynamics.
The non-Markovian terms are essential for accurately capturing these small-scale patterns before a rare or typical event occurs. When large-scale mode amplitudes are near zero during interval I (Table 2), the optimal parameterization (Eq. (86)) is dominated by the (optimized) non-Markovian field $\xi_{t}$ given by:
@@
$\hskip 142.26378pt+Z^{n}_{\tau_{n}^{\ast}}(t;\omega).$
After this initial transient period, nonlinear effects become significant, and the large-scale mode amplitudes undergo substantial fluctuations. Panels C-F in Figures 5 and 6 demonstrate this behavior and the optimal reduced model’s ability to predict these large-scale fluctuations beyond interval I. The OPM reduced system’s ability to capture such transitions is rooted in its very structure. Its coefficients in Eq. (85) are nonlinear functionals of the non-Markovian $M_{n}$-terns within $\xi_{t}$ allowing for an accurate representation of the genuine nonlinear interactions between noise and nonlinear terms in the original sACE, which drive fluctuations in the large-mode amplitudes (Modes 1 to 4).
After a transition occurs and energy is transferred to the low modes, the stochastic component $\xi_{t}$ and its non-Markovian terms become secondary while the nonlinear terms in the stochastic parameterization (82) become crucial for capturing the average behavior of the sACE dynamics, as highlighted in Section V.5 below.
@@ V.4 Rare and typical events statistics from non-Markovian reduced models
Thus, the optimal reduced system reproduces faithfully the probability distribution of both common (frequently occurring sign-change profiles) and rare (constant-sign profiles) final states at time $T$ (Fig. 7A). This is particularly impressive considering the model was trained on just a single path, yet generalizes well to predict the behavior across a large ensemble of out-of-sample paths.
Figure 8A shows the training stage over a typical transition path. The optimization of the free parameter $\tau$ in the stochastic parameterization $\Phi_{n}$ (see Eq. (82)) is executed over a typical transition path, by minimizing the normalized parameterization defects $Q_{n}(\tau,T)=\mathcal{Q}_{n}(\tau,T)/\overline{\big{|}u_{n}(t)\big{|}^{2}}$ ($\mathcal{Q}_{n}(\tau,T)$ defined in Eq. (53)) for all the relevant $n$ (here $n=5,\cdots,8$). We observe that $Q_{5}$ and $Q_{6}$ exhibit clear minima, whereas the other modes have their parameterization defect not substantially diminished at finite $\tau$-values compared to its asymptotic value (as $\tau\rightarrow\infty$). Same features are observed over a rare transition path; see Figure 8B. Recall that our reduced system is an ODE system whose stochasticity comes exclusively from its random coefficients as the resolved modes in the full system are not directly forced by noise. It is this specificity that enables the reduced equations to effectively predict noise-induced transitions, including rare events. This remarkable capability relies on the optimized memory content embedded within the path-dependent, non-Markovian coefficients.
Thus, the optimized, non-Markovian and nonlinear reduced equation (85) is able to track efficiently how the noise, acting on the “unseen” part of the system, interacts with the resolved modes through these non-Markovian coefficients. While these coefficients effectively encode the fluctuations for a good reduction of the sACE system, they alone are insufficient to handle regimes with weak time-scale separation such as dealt with here. Understanding the average behavior of the resolved dynamics is equally important. The next section explores how our parameterization, particularly its nonlinear terms, play a crucial role in capturing this average behavior.
@@ V.5 Weak time-scale separation, nonlinear terms and conditional expectation
We assess the ability of our reduced system to capture this average dynamical behavior in the physical space. In that respect, Figures 7B and 7D show the expected profiles of the sACE’s states obtained at $t=T$ (solid black curves) from the sACE (76), along with their standard deviation (grey shaded areas), for the ensemble used for Fig. 7A. Similarly, Figures 7C and 7E show the expected profiles (solid red curves) with their standard deviation as obtained at $t=T$ from $u^{\mathrm{opm}}(T,x,\cdot)$ defined in Eq. (87), once the low-mode amplitude is simulated by the optimal reduced model (86) and lifted to $H_{\mathfrak{f}}$ by using the stochastic parameterization $\Phi_{{\bm{\tau}}^{\ast}}$.
@@
Figure 9: Markovian optimal parameterization for the 5th and 6th modes’ amplitudes. Panel A: The black curve shows the time evolution of the 6th mode’s amplitude, $u_{6}(t)$, obtained by solving the governing equation, Eq. (76), over the time interval 0 to 400. This 6th mode’s amplitude is shown against the corresponding first and second modes’ amplitudes, $u_{1}(t)$ and $u_{2}(t)$. For visualization purposes, the other solution components in the four-dimensional subspace $H_{\mathfrak{c}}$, $u_{3}(t)$ and $u_{4}(t)$, are fixed at their most probable values, $p_{3}$ and $p_{4}$, respectively. The manifold shown is the underlying Markovian OPM for the 6th mode’s amplitude (Eq. (89) with $n=6$). Panel B: Same for the 5th mode’s amplitude $u_{5}(t)=\langle u(t),\bm{e}_{5}\rangle$ and its corresponding Markovian OPM (Eq. (89) with $n=5$).
@@
By conducting large-ensemble (online) simulations of our reduced model (Eq. (86)), we accurately reproduce the mean motion and second-order statistics of the original sACE. This success stems from our parameterization’s ability to effectively represent the average behavior of the unresolved variables during the offline training phase, as defined by the probability measure $\mu$ in (58). Figure 9 illustrates this point for the fifth and sixth modes’ amplitudes of the sACE system. Notably, the nonlinear nature of the Markovian optimal parameterization, evident after taking the expectation (see Eq. (89) below), highlights the crucial role of nonlinear terms in our stochastic parameterization (Eq. (82)) in achieving this accurate representation.
The manifold shown in Fig. 9A (resp. Fig. 9B) is obtained as the graph of the following Markovian OPM for the 6th (resp. 5th) mode’s amplitude, given by the expectation of Eq. (82) with $n=6$ (resp. $n=5$) after optimization (Fig. 8),
$(X_{1},X_{2})\mapsto$ $\mathbb{E}[\Phi_{n}(\tau_{n}^{\ast},\bm{X})]$ (89)
$=\sum_{i,j,k=1}^{q}\Big{(}E_{ijk}^{n}(\tau_{n}^{*})C_{ijk}^{n}X_{i}X_{j}X_{k}\Big{)},$
for $\bm{X}=(X_{1},X_{2},X_{3},X_{4})$ with $X_{3}=p_{3}$ and $X_{4}=p_{4}$, where $p_{3}$ and $p_{4}$ denote the most probable values of the 3rd and 4th solution components, $u_{3}(t)$ and $u_{4}(t)$, to enable visualization as a 2D surface while favoring a certain ”typicalness” of the visualized manifold. Note that the Markovian OPM corresponds simply to the deterministic cubic term in (82) after replacing $\tau$ by the optimal $\tau_{n}^{*}$, since the remaining stochastic terms equal to $\sigma_{n}\int_{t-\tau_{n}^{*}}^{t}e^{\lambda_{n}(t-s)}\,\mathrm{d}W^{n}_{t}(\omega)$ (see Remark III.1), whose expectation is zero as stochastic integral in the sense of Itô.
@@
Without the optimization stage, the parameterizations of Section II experience deficient accuracy when the memory content is taken to be $\tau=\infty$, due to small values that the spectral gaps $\delta_{ijk}^{n}$ take in denominators. As a reminder, small spectral gaps are indeed a known limitation of traditional invariant manifold techniques, often leading to inaccurate results in this case. Optimizing the parameter $\tau$ helps address this issue. It introduces corrective terms like $(1-\exp(-\delta_{ijk}^{n})\tau)$ in the cubic coefficients ($E_{ijk}^{n}$, Eq. (44)) that effectively compensate for these small gaps. This is especially important for fifth and sixth (unresolved) modes whose parameterization defects show a clear minimum during optimization (Figure 8).
@@
Thus, the closure results shown in Figure 7 demonstrate our reduction approach’s effectiveness in handling weak time-scale separation scenarios, while also suggesting potential for further improvement (see Section IX). Our data-driven optimization of analytical parameterization formulas offers a compelling alternative to traditional multiscale methods or those based on invariant manifolds [132, 103, 102, 133, 134]. These methods, while providing rigorous error estimates, often require strong time-scale separation for the derivation of their amplitude equations. Additionally, they can achieve higher-order nonlinearities in their reduced equations too but through a different route: the Itô formula is applied after exploiting a slow timescale of the system (e.g., [103, Eq. (12)] and [136, Eq. (15)]). Integrating optimization concepts into these alternative frameworks could be a promising avenue to explore how it performs in weak time-scale separation regimes.